Please help on this finance!

Griffey Communications recently realized $122,500 in operating income. The company had interest income of $25,000 and realized $70,000 in dividend income. The company's interest expense was $40,000. Using the corporate tax schedule below, what is Griffey's tax liability?

Taxable Income Tax on Base of Bracket Percentage on Excess above Base
Up to $50,000 $0 15%
$50,000-$75,000 7,500 25   
$75,000-$100,000 13,750 34   
$100,000-$335,000 22,250 39   
$335,000-$10,000,000 113,900 34   
$10,000,000-$15,000,000 3,400,000 35   
$15,000,000-$18,333,333 5,150,000 38   
Over $18,333,333 6,416,667 35   

Assume a 70% dividend exclusion for tax on dividends.

a. $27,359
b. $36,702
c. $37,702
d. $33,365
e. $32,364
Answer #1

Computation of gross total income of the company:

Operating income : 122500

+ interest income ; 25000

+ dividend income : 21000(70% of dividend is excluded for tax purposes)

-interest expense : 40000

Total = 128500

Tax base as per the rates table up to the amount 100000 = 22250

tax on excess 28500 at 39% = 11115

So, the total tax liability = 22250+11115 = 33365(D)

So option = D